OpenStudy (anonymous):

What equation can be used to calculate the time of impact when you have the impulse and force applied to an object?

OpenStudy (doc.brown):

Impulse is force times time. J=Ft

OpenStudy (anonymous):

thats not an option. thats what i thought too. here are the options; @doc.brown

OpenStudy (doc.brown):

In their notation\[I=F(\Delta t)\] Does that help?

OpenStudy (anonymous):

so to rearrange it would it just be t=f/i or i/f? @doc.brown

OpenStudy (doc.brown):

How do you get \(\Delta t\) by itself?

OpenStudy (anonymous):

you would add it to the other side and cancel out, correct? @doc.brown

OpenStudy (doc.brown):

Divide both sides by F.

OpenStudy (anonymous):

meaning it's \[\Delta T = \frac{ I}{ F}\] @doc.brown ?

OpenStudy (doc.brown):

Yup

OpenStudy (doc.brown):

\[I=F(\Delta t)\]\[F(\Delta t)=I\]\[\frac{F(\Delta t)}{F}=\frac{I}{F}\]\[\frac{\cancel{F}(\Delta t)}{\cancel{F}}=\frac{I}{F}\]\[\Delta t=\frac{I}{F}\]